OpenStudy (anonymous):

Putting this into point/slope form using the equation below. Points given are: (4, -6) and (4, 3).

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

okay here u have to use the formula (y−y1)/(y2−y1)=(x−x1)/(x2−x1) where x1,y1 is one point and x2,y2 is the 2nd point!!

OpenStudy (anonymous):

I just gave the formula...

OpenStudy (anonymous):

yeah but using this formula its easier!! anyway the answer is y=-6

OpenStudy (anonymous):

Wouldn't it be undefined though? Because when finding the slope, you would subtract the two y's and get -9, and subtract the two x's and get 0, thus making it undefined on a graph.

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

tan a=slope if slope is undefined then a=90 degress. therefore the solution is a st line perpendicular to the x axis.

OpenStudy (anonymous):